Two resistors (R1 =3Ω&R2 =2Ω) are wired in series in a DC circuit. What is the equivalent resistance?
Answer #1

As for series combination of resistances

Equivalent resistance, R = R1 + R2 = 3 + 2 = 5 ohm.